OpenStudy (anonymous):

What is the slope of the line between (-4, 4) and (-1, -2)? 1 2 -2 -1

hartnn (hartnn):

The slope of the line through points (x1,y1) and (x2,y2) is given by : \(\huge m=\frac{y_1-y_2}{x_1-x_2}\) now,just put the values and find m.

OpenStudy (anonymous):

Thanks that helps!

hartnn (hartnn):

welcome ^_^

hartnn (hartnn):

what slope did you get?? just to verify...

OpenStudy (anonymous):

umm i dont think i did it right...i got -0.5

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

yeah, m = 4-(-2) / -4- (-1) = 6/ -3 = -2

OpenStudy (anonymous):

ohh i did it backwards..oops lol but thankss

hartnn (hartnn):

welcome :)